OpenStudy (anonymous):

Write an equation of the line perpendicular to the line given by 2x + 3y = 7 and having the same y-intercept as the line given by 5x + 2y = 10

OpenStudy (precal):

solve both of them for y

OpenStudy (anonymous):

i got 3x - 2y = 7 is perpendicular to 2x + 3y = 7 the y-intercept (when x=0) of 5x + 2y = 10 is 5 , coordinates (0,5) plug the coordinates into the new equation to get the proper y-intercept 3(0) - 2(5) = -10 the correct equation is 3x - 2y = -10

OpenStudy (precal):

 slope is -2/3